# Sample and Map from a Single Convex Potential: Generation using Conjugate Moment Measures

### v1

@@ A.5 Hyperparameters and infrastructure
The hyperparameters used for the MNIST experiments are given in Table 2. To determine these hyperparameters, we used commonly applied architectures for ICNNs and conducted a grid search over $\gamma$ testing values 1e-3, 5e-4, 1e-4, 5e-5 and 1e-5, as well as for the learning rate of $w_{\theta}$, considering values 1e-3, 5e-4 and 1e-4. We adopt the approach of Amos (2023), where the cost of computing the conjugate at each training step is amortized by initializing the conjugate solver with the predictions of an MLP. This MLP is trained via regression on the outputs of the conjugate solver. The MLP consists of three hidden layers with 128 units each for the 2D experiments and three hidden layers with 256 units each for the MNIST experiment. It is trained using the Adam optimizer with default parameters, and a fixed learning rate of 1e-4 for the 2D experiments and 5e-4 for the MNIST experiment. It is important to note that the number of LMC steps listed in the tables refers to the steps taken to sample from $\mathfrak{P}_{w_{\theta}}$ starting from uniform noise, as used to generate the figures. During training, however, we reuse the particles sampled from the previous gradient step and apply $200$ LMC steps to these particles to form the new batch.
@@
Conjugate solver $w^{*}_{\theta}$ Adam
step size with cosine decay schedule
init value = 0.1
alpha = 1e-4
$\beta_{1}=0.9$
$\beta_{2}=0.999$
100 iterations
Langevin sampling number of steps $=6000$ (=$200$ during training)
number of gradient steps 700,000
Table 1: Hyperparameters for the 2D experiments of Figure 3.
@@
Langevin sampling step size $\gamma=0.001$
number of steps $=6000$ (=$200$ during training)
@@
batch size 512
Table 2: Hyperparameters for the MNIST experiment of Figure 4.

### v2

@@ E.1 Hyperparameters for CMFGen
Our algorithm, CMFGen, introduces no additional hyperparameters compared to the generative ICNN trained with the solver from Amos (2023), except for the number of LMC steps used when sampling from $\mathfrak{P}_{w_{\theta}}$. For a fair comparison, we adopt the same ICNN architecture for both CMFGen and the generative ICNN baseline shown in Figures 13, 5, and 15. We follow the standard ICNN design from Vesseron and Cuturi (2024), which is known to perform well in generative settings. All the 2D experiments in Figure 4 were generated using a common set of hyperparameters, detailed in Table 1. Only the number of particles used (i.e., batch size) varies across experiments. For the 2D distributions Circles, S-curve, Scaled-Rotated S-curve and Diag-Checkerboard we used $1024$ particles, while for Checkerboard, a larger number of particles ($4096$) was used to ensure the stability of the algorithm. The hyperparameters used for the MNIST and Cartoon experiments are given in Table 2. Moreover, we adopt the approach of Amos (2023), where the cost of computing the conjugate at each training step is amortized by initializing the conjugate solver with the predictions of an MLP. This MLP is trained via regression on the outputs of the conjugate solver. The MLP consists of three hidden layers with 128 units each for the 2D experiments and three hidden layers with 256 units each for the MNIST and Cartoon experiment. It is trained using the Adam optimizer with default parameters, and a fixed learning rate of 1e-4 for the 2D experiments and 5e-4 for the MNIST experiment. As explained in Amos (2023), the predictions of the MLP are then refined with a conjugate solver whose hyperparameters are given Table 3 and are kept consistent across all experiments with both CMFGen and the generative ICNN. It is important to note that the number of LMC steps listed in the tables refers to the steps taken to sample from $\mathfrak{P}_{w_{\theta}}$ starting from uniform noise, as used to generate the figures. During training, however, we reuse the particles sampled from the previous gradient step and apply $200$ LMC steps to these particles to form the new batch. The number of LMC steps during training reflects a trade-off between computational efficiency and performance: using significantly fewer steps (e.g., $50$) results in degraded performance.
@@
Langevin sampling number of steps $=10000$ (=$200$ during training)
number of gradient steps 1,000,000
Table 1: Hyperparameters for CMFGen and the generative ICNN in 2D experiments. The only additional hyperparameter used by CMFGen, compared to the generative ICNN, is denoted in orange.
@@
Langevin sampling number of steps $=10000$ (=$200$ during training)
number of gradient steps 50,000
batch size 512
Table 2: Hyperparameters for CMFGen and the generative ICNN in the MNIST and Cartoon experiments. The only additional hyperparameter used by CMFGen, compared to the generative ICNN, is denoted in orange.
Hyperparameter Value
@@
Table 3: Hyperparameters for the conjugate solver used in CMFGen and for the generative ICNN (the same hyperparameters are used in all experiments).
E.2 Hyperparameters for CMFMA
The hyperparameters for our CMFMA algorithm are listed in Tables 4 and 5. We follow standard ICNN architectures, except for the 1D experiments, where we observed discontinuities in the learned potential when using ELU activations. This issue arises because the second derivative of ELU is not continuous, and the CMFMA regression loss involves the Hessian of the network $w_{\theta}$. To address this, we use Softplus activations with a beta coefficient of $10$.
Hyperparameter Value
$w_{\theta}$ architecture dense
Softplus($\beta=10.0$) activation functions
$w_{\theta}$ optimizer Adam
step size $=0.0001$
$\beta_{1}=0.5$
$\beta_{2}=0.5$
@@ E.2 Hyperparameters for CMFMA
batch size 1024
Table 4: Hyperparameters for CMFMA in 1D experiments.
Hyperparameter Value
$w_{\theta}$ architecture dense
ELU activation functions
$w_{\theta}$ optimizer Adam
step size $=0.0001$
$\beta_{1}=0.5$
$\beta_{2}=0.5$
number of gradient steps 500,000
batch size 1024
Table 5: Hyperparameters for CMFMA in 2D experiments.